# A Nonperturbative Toolkit for Quantum Gravity

### v2

@@ 6.4 Other possible bases and saturation of the spectral form factor
By the argument in Sec. 5, any set of $\kappa$ states that satisfies $\overline{\braket{i}{j}}\propto\delta_{ij}$ with $\overline{\braket{i}{j}\braket{j}{i}}|\neq 0$ for any two states $|i\rangle,|j\rangle$ in the set should form a basis as $\kappa\to\infty$. The shell states are useful because $\overline{\braket{i}{j}\braket{j}{i}}\neq 0$ already in the semiclassical saddlepoint approximation, and corresponding wormholes are easily constructed. A potential alternative basis for $\mathcal{H}_{LR}$ is the time-evolved TFD states recently discussed in Banerjee:2024fmh . This family of states is obtained by evolving the state $|\beta\rangle$ by a varying amount in Lorentzian boundary time. As discussed in Magan:2024nkr , proving that these states span the full quantum gravity Hilbert space implies the saturation of the plateau of the gravity spectral form factor. In chaotic quantum systems, time evolved TFD states satisfy the two above properties as shown in Verlinde:2021kgt . If this can be shown explicitly for the gravity construction of these states, a corollary is that the spectral form factor for gravity in any dimension has a late time plateau. However, the relevant gravitational wormholes have not been constructed, and may not exist as saddlepoints, because of the absence of matter stress energy supporting the wormhole topology.
@@
One way of achieving greater control is to create a bubble of Euclidean de Sitter within Euclidean asymptotically anti-de Sitter spacetime. To achieve this, Balasubramanian:2023xyd imposed an additional constraint on the Euclidean AdS path integral that the relevant saddlepoints must contain an asymptotic space-like boundary at future infinity after continuation to Lorentzian signature. This procedure can be seen as imposing a mixed Euclidean/Lorentzian boundary condition that requires the saddle geometries to be locally dS in the infinite future. This rule could be used in conjunction with the highly simplified AdS path integral setting provided by our toolkit to probe de Sitter quantum gravity.

### v3

@@ 6.4 Other possible bases and saturation of the spectral form factor
By the argument in Sec. 5, any set of $\kappa$ states that satisfies $\overline{\braket{i|j}}\propto\delta_{ij}$ with $\overline{\braket{i|j}\braket{j|i}}|\neq 0$ for any two states $|i\rangle,|j\rangle$ in the set should form a basis as $\kappa\to\infty$. The shell states are useful because $\overline{\braket{i|j}\braket{j|i}}\neq 0$ already in the semiclassical saddlepoint approximation, and corresponding wormholes are easily constructed. A potential alternative basis for $\mathcal{H}_{LR}$ is the time-evolved TFD states recently discussed in Banerjee et al. (2024). This family of states is obtained by evolving the state $|\beta\rangle$ by a varying amount in Lorentzian boundary time. As discussed in Magan and Wu (2024), proving that these states span the full quantum gravity Hilbert space implies the saturation of the plateau of the gravity spectral form factor. In chaotic quantum systems, time evolved TFD states satisfy the two above properties as shown in Verlinde (2021). If this can be shown explicitly for the gravity construction of these states, a corollary is that the spectral form factor for gravity in any dimension has a late time plateau. However, the relevant gravitational wormholes have not been constructed, and may not exist as saddlepoints, because of the absence of matter stress energy supporting the wormhole topology.
@@
One way of achieving greater control is to create a bubble of Euclidean de Sitter within Euclidean asymptotically anti-de Sitter spacetime. To achieve this, Balasubramanian et al. (2024e) imposed an additional constraint on the Euclidean AdS path integral that the relevant saddlepoints must contain an asymptotic space-like boundary at future infinity after continuation to Lorentzian signature. This procedure can be seen as imposing a mixed Euclidean/Lorentzian boundary condition that requires the saddle geometries to be locally dS in the infinite future. This rule could be used in conjunction with the highly simplified AdS path integral setting provided by our toolkit to probe de Sitter quantum gravity.